# Theory Uncertainties in the Extraction of $\alpha_{s}$ from Drell-Yan at Small Transverse Momentum

### v1

@@ 4.2.4 Dependence on pre-fit theory constraints
Figure -952: Post-fit constraints on TNPs at N3+1LL profiled against the N4LL central prediction using different pre-fit theory constraints, left panel: $\Delta\theta_{n}=1$, central panel: $\Delta\theta_{n}=2$ and right panel: $\Delta\theta_{n}=4$. The stars indicate the true values of the TNPs.
@@
In figure -952, we show the resulting pull plots. As the pre-fit uncertainty on the TNPs is increased, also the post-fit uncertainty on the TNPs increases but much less so for some TNPs. Overall, the post-fit values become more and more compatible within uncertainties with the TNPs’ true values, because it becomes easier for the fit to move the TNPs to their true values. With $\Delta\theta_{n}=4$, the constraints on certain TNPs become stronger relative to the pre-fit constraint and the less constrained TNPs. This is expected since with $\Delta\theta_{n}=4$ the TNPs become essentially unconstrained, which makes the differing sensitivities to the various TNPs more clear as the data acts to constrain the most relevant TNPs. When relaxing the theory constraint, the TNPs having the strongest impact on $\alpha_{s}(m_{Z})$ remain the same, namely $\gamma_{\nu}$, $S$, and $B_{qg}$. However, their individual impacts reduce while the impact of the other TNPs increases.
@@
Figure -951: Perturbative uncertainties on $\alpha_{s}(m_{Z})$ when profiling TNPs at N3+1LL against the N4LL central prediction using different pre-fit theory constraints $\Delta\theta_{n}=1,2,4$. The displayed uncertainties are in units of $10^{-3}$.
@@
In figure -951, we show the results for $\alpha_{s}(m_{Z})$ for all three cases. Interestingly, despite the shifts in the impacts of individual TNPs, the total uncertainty on $\alpha_{s}(m_{Z})$ remains essentially unchanged. This shows that the total uncertainty after profiling is essentially driven by the constraints imposed by the data. The precise theory constraint does not affect the total uncertainty but only the relative distribution of the uncertainties among the TNPs. This is the case in a situation like ours here where the uncertainties of the data are $\lesssim$ the pre-fit theory uncertainties, i.e., in a situation where without profiling we would be limited by theory uncertainties. In the opposite situation, where the data uncertainties are larger than the pre-fit theory uncertainties, the theory uncertainty will depend on the theory constraint also after profiling simply because the data will be too weak to constrain the theory uncertainties. However, in this situation the theory uncertainties are subdominant to begin with, and whilst the theory uncertainty itself will depend more on the adopted theory constraint, the total uncertainty will not. Therefore, as discussed before, the only situation to be avoided is that the inherent precision of the theory model, i.e., the order to which TNPs are included is too low compared to the data precision such that it becomes overconstrained. In this case, the theory model can be improved by including additional TNP orders.
@@
As already discussed in the context of figure -953, there is a minor bias in the central value of $\alpha_{s}(m_{Z})$ due to the $\Delta\theta_{n}=1$ theory constraint. Now, as the theory constraint is weakened, the true values of the TNPs lie well within it, so the associated penalties are reduced. As a result, the bias further reduces for $\Delta\theta_{n}=2$ and disappears completely for $\Delta\theta_{n}=4$.
@@ 5 Nonperturbative Effects
Our nominal result thus corresponds to profiling the TNPs at N3+1LL along with $\lambda_{2,4}$ for the CS kernel and $\Lambda_{2,4}$ for the TMD PDFs. For $\lambda_{2,4}$ we impose the representative lattice results discussed in section 3.3.2 as prior Gaussian constraints. As we will see, including the lattice constraints for $\lambda_{2,4}$ helps guide the exploration of parameter space, allowing the data to determine the TMD parameters $\Lambda_{2,4}$, which contribute to the uncertainty but for which we otherwise lack constraints from first-principles calculations.
@@
Figure -950: Left panel: Uncertainties in the $q_{T}$ spectrum at N3+1LL relative to the N4LL, due to the nonperturbative parameters (purple band), the TNPs (orange band) and their sum in quadrature band (in dark orange). The different lines show the post-fit relative impact of each nonperturbative parameter in the spectrum. Right panel: Post-fit constraints on the TNPs and nonperturbative parameters (error bars) and their impact on $\alpha_{s}(m_{Z})$, with the solid (dashed) grey band showing the impact of the post-fit downward (upward) variations of each parameter. The stars indicate the true value of the parameters. The post-fit constraints for $\Lambda_{2}$ and $\lambda_{2}$ ($\Lambda_{4}$ and $\lambda_{4}$) are normalized to $0.25\,\mathrm{GeV}^{2}$ ($0.06\,\mathrm{GeV}^{4}$).
@@
In figure -950, we show the result of our nominal fit including nonperturbative parameters. The left panel displays the breakdown of the post-fit uncertainties in the $q_{T}$ spectrum. The uncertainty due to the nonperturbative parameters is shown in violet, the perturbative uncertainty due to TNPs is shown in orange, and their combined uncertainty is represented by the dark orange band, which corresponds to the total post-fit uncertainty for this Asimov test. [^7] When profiling multiple sources of uncertainty, the breakdown of the total uncertainty into individual components is somewhat ambiguous. Here, we obtain the individual bands on the $q_{T}$ spectrum by considering the relevant sub-matrices of the total post-fit covariance matrix. One benefit of the Asimov approach however is that we can also consider different sources in isolation, as was done for the TNP component in section 4.2. We do not show a pre-fit band here, since we do not have pre-fit constraints for $\Lambda_{2,4}$. The perturbative pre-fit uncertainty corresponds to the one shown in figure -954. Overlaid on the plot are the post-fit relative impacts of each nonperturbative parameter. A strong negative correlation emerges between $\lambda_{2}$ and $\Lambda_{2}$, as evident from their significant impact on the spectrum despite the relatively small size of the overall nonperturbative uncertainty. In the right panel of the same figure, we show the constraints on all fit parameters, along with their individual impacts on $\alpha_{s}(m_{Z})$. Compared to the purely perturbative case discussed in section 4.2.3, the post-fit constraints on the TNPs are still consistent with their true values (indicated by stars), but some are now not as strongly constrained by the data. Due to the additional fit parameters introduced, the data has additional freedom now to determine which parameters are most relevant in reducing the uncertainty on the parameter of interest. It is clear that some nonperturbative parameters, especially $\Lambda_{2,4}$, have a large impact on $\alpha_{s}(m_{Z})$.
@@
Figure -949: Uncertainty on $\alpha_{s}(m_{Z})$ from various Asimov tests. Red bars include TNP profiling (i.e., fitting $\alpha_{s}(m_{Z})$ together with all theory nuisance parameters). Below, we show cases excluding TNP profiling, in blue, with different sets of fitted parameters. From top to bottom: fit only $\alpha_{s}(m_{Z})$, fit $\alpha_{s}(m_{Z})$, $\lambda_{2,4}$, $\Lambda_{2,4}$ including lattice QCD constraints for CS kernel parameters, fit $\alpha_{s}(m_{Z})$, $\lambda_{2}$ and $\Lambda_{2,4}$, fit $\alpha_{s}(m_{Z})$ and $\Lambda_{2,4}$. All the uncertainties are in units of $10^{-3}$.
@@
Next, we explore the relationship between the nonperturbative effects and the uncertainty on $\alpha_{s}(m_{Z})$, and how the TNPs behave in this context. Figure -949 presents the results of several Asimov fits as discussed next. For each case, we display two uncertainty bars. The outer red bars show the total uncertainty of the nominal fit simultaneously profiling TNPs and nonperturbative parameters. The inner blue bars show the uncertainty associated with fitting $\alpha_{s}(m_{Z})$ alongside the nonperturbative parameters only. They are obtained by repeating the fit fixing the TNPs to their nominal post-fit central values.
@@
Fortunately, as discussed in section 3.3, the CS kernel can be constrained using lattice QCD data. Incorporating these constraints – via both the central values from Eq. (3.3.2) and the associated covariance matrix in Eq. (3.35) – enables us to include both $\lambda_{2,4}$ in the fit, alongside the TMD PDF parameters $\Lambda_{2,4}$. This yields our nominal fit already shown in figure -950. The $\alpha_{s}(m_{Z})$ uncertainties resulting from this fit are
@@ Appendix A Comments on Recent $\alpha_{s}$ Extraction by ATLAS
In table 2, we summarize our estimates of the expected size of perturbative uncertainties based on fits to Asimov data and compare them to the values quoted in Ref. [33]. For scale variations, we use the quadrature sum of envelopes here. Using instead the naive total envelope ($1.73\times 10^{-3}$) leads to a slightly smaller total perturbative uncertainty of ±$\pm 2.30\times 10^{-3}$. Our uncertainty estimates are based on fits to Asimov data, so while the precise numbers will differ in a fit to the real data their typical size will be the same. For the uncertainties due to scale variations, we have explicitly checked in preliminary fits to real data that the size of the obtained uncertainty is indeed very similar. We thus have reason to believe that the total theory uncertainty quoted in Ref. [33] is underestimated by a factor of 4 to 5. This would make it by far the dominant uncertainty and as a result also the total uncertainty of ±$\pm 0.9\times 10^{-3}$ quoted in Ref. [33] would be underestimated by about a factor of 3.
@@
Scale variations ±$\pm 0.42$$\pm 2.43$
N4LL approximation ±$\pm 0.04$$\pm 0.75$
Flavor/quark masses $+0.40$ $-0.29$$\pm 1.32$
Total $+0.58$ $-0.51$$\pm 2.87$
@@ A.1 Perturbative uncertainties from scale variations
The analysis in Ref. [33] uses the method of enveloped scale variations discussed in section 2.2 to estimate the perturbative theory uncertainty on the resummed $q_{T}$ spectrum. This uncertainty is propagated to the fitted $\alpha_{s}(m_{Z})$ by scanning over individual scale variations. This yields maximum deviations in $\alpha_{s}(m_{Z})$ of $+0.23\times 10^{-3}$ and $-0.61\times 10^{-3}$, which are symmetrized to the quoted ±$\pm 0.42\times 10^{-3}$ scale-variation uncertainty on $\alpha_{s}(m_{Z})$.
@@
As we have argued in the main text, extracting $\alpha_{s}(m_{Z})$ from the small-$q_{T}$ spectrum crucially relies on shape effects and is therefore extremely sensitive to the theory uncertainty on the shape of the spectrum and the corresponding bin-by-bin theory correlations. As discussed in detail in section 2.2, these are not correctly captured by scale variations, which means the resulting variations in the fitted $\alpha_{s}(m_{Z})$ cannot be regarded as a trustworthy uncertainty estimate. Scanning over scale variations in section 4.1 we find significantly larger maximum variations in the fitted $\alpha_{s}(m_{Z})$, as seen in the right panel of figure -958. The envelope over all scale variations yields $1.73\times 10^{-3}$, while adding in quadrature the envelopes of different classes of scale variations yields $2.43\times 10^{-3}$, see Eqs. (4.1) and (4.1).
[…]
@@
One difference which could (at least partially) contribute to this large discrepancy is that the scale-variation recipes used in the SCETlib and DYturbo resummation frameworks differ. Whilst both use the conventional factor of two for varying scales up and down and avoid simultaneous variations that would compound to factors of four in any scale ratios, the specific types of scale variations differ substantially and are more limited in DYturbo. In particular, the low scale $\sim q_{T}\sim 1/b_{T}$ is varied as part of the scale variation setup in SCETlib, while it is kept fixed in DYturbo. We stress that there are no compelling theoretical arguments for keeping this low scale fixed when estimating uncertainties using scale variations. [^11] This low scale is genuine to the resummation problem and therefore must effectively appear in any resummation framework. It is also kept variable in the original CSS framework [69] on which DYturbo is based. In the context of an ongoing benchmarking study of various $q_{T}$ resummation frameworks within the LHC electroweak working group, the need for varying this low scale has been widely recognized by now and, as far as we can tell, has been gradually adopted also by most other resummation frameworks.
@@
As discussed in section 2, in a situation like this where theory correlations are critically important, for the same overall uncertainty in the spectrum, the resulting uncertainty on $\alpha_{s}(m_{Z})$ entirely depends on the specific scale variations that happen to be included in the scanning. A similar breakdown into individual scale variations as in our figure -958 of the results in Ref. [33] might shed some more light on this. Ultimately, however, the large discrepancies we observe between different resummation frameworks using different scale variation recipes likely just reflect the fact that the scale-variation method cannot be relied upon for propagating the perturbative uncertainty from the $q_{T}$ spectrum to $\alpha_{s}(m_{Z})$.
@@
Nevertheless, if scale variations are the method of choice, one must be aware of their limitations and at minimum it must be ensured that a given scale-variation recipe does not underestimate the resulting uncertainty on the parameter of interest. For this purpose, we have clear evidence that the type of scale variations performed in SCETlib yield a more realistic uncertainty for $\alpha_{s}(m_{Z})$. This is because when scanning over TNPs, which do encode the correct theory correlations, spectrum uncertainties of comparable (or even somewhat smaller) size translate into an uncertainty on $\alpha_{s}(m_{Z})$ of $1.75\times 10^{-3}$ (see section 4.2.1 and figure -957), which is similar to what we find with scale variations.
@@
A.2 N4LL approximation uncertainties

### v2

@@ 4.2.4 Dependence on pre-fit theory constraints
Figure 8: Post-fit constraints on TNPs at N3+1LL profiled against the N4LL central prediction using different pre-fit theory constraints, left panel: $\Delta\theta_{n}=1$, central panel: $\Delta\theta_{n}=2$ and right panel: $\Delta\theta_{n}=4$. The stars indicate the true values of the TNPs.
@@
In figure 8, we show the resulting pull plots. As the pre-fit uncertainty on the TNPs is increased, also the post-fit uncertainty on the TNPs increases but much less so for some TNPs. Overall, the post-fit values become more and more compatible within uncertainties with the TNPs’ true values, because it becomes easier for the fit to move the TNPs to their true values. With $\Delta\theta_{n}=4$, the constraints on certain TNPs become stronger relative to the pre-fit constraint and the less constrained TNPs. This is expected since with $\Delta\theta_{n}=4$ the TNPs become essentially unconstrained, which makes the differing sensitivities to the various TNPs more clear as the data acts to constrain the most relevant TNPs. When relaxing the theory constraint, the TNPs having the strongest impact on $\alpha_{s}(m_{Z})$ remain the same, namely $\gamma_{\nu}$, $S$, and $B_{qg}$. However, their individual impacts reduce while the impact of the other TNPs increases.
@@
Figure 9: Perturbative uncertainties on $\alpha_{s}(m_{Z})$ when profiling TNPs at N3+1LL against the N4LL central prediction using different pre-fit theory constraints $\Delta\theta_{n}=1,2,4$. The displayed uncertainties are in units of $10^{-3}$.
@@
In figure 9, we show the results for $\alpha_{s}(m_{Z})$ for all three cases. Interestingly, despite the shifts in the impacts of individual TNPs, the total uncertainty on $\alpha_{s}(m_{Z})$ remains essentially unchanged. This shows that the total uncertainty after profiling is essentially driven by the constraints imposed by the data. The precise theory constraint does not affect the total uncertainty but only the relative distribution of the uncertainties among the TNPs. This is the case in a situation like ours here where the uncertainties of the data are $\lesssim$ the pre-fit theory uncertainties, i.e., in a situation where without profiling we would be limited by theory uncertainties. In the opposite situation, where the data uncertainties are larger than the pre-fit theory uncertainties, the theory uncertainty will depend on the theory constraint also after profiling simply because the data will be too weak to constrain the theory uncertainties. However, in this situation the theory uncertainties are subdominant to begin with, and whilst the theory uncertainty itself will depend more on the adopted theory constraint, the total uncertainty will not. Therefore, as discussed before, the only situation to be avoided is that the inherent precision of the theory model, i.e., the order to which TNPs are included is too low compared to the data precision such that it becomes overconstrained. In this case, the theory model can be improved by including additional TNP orders.
@@
As already discussed in the context of figure 7, there is a minor bias in the central value of $\alpha_{s}(m_{Z})$ due to the $\Delta\theta_{n}=1$ theory constraint. Now, as the theory constraint is weakened, the true values of the TNPs lie well within it, so the associated penalties are reduced. As a result, the bias further reduces for $\Delta\theta_{n}=2$ and disappears completely for $\Delta\theta_{n}=4$.
@@ 5 Nonperturbative Effects
Our nominal result thus corresponds to profiling the TNPs at N3+1LL along with $\lambda_{2,4}$ for the CS kernel and $\Lambda_{2,4}$ for the TMD PDFs. For $\lambda_{2,4}$ we impose the representative lattice constraints in Eqs. (3.3.2) and (3.35) as prior Gaussian constraints. As we will see, including the lattice constraints for $\lambda_{2,4}$ helps guide the exploration of parameter space, allowing the data to determine the TMD parameters $\Lambda_{2,4}$, which contribute to the uncertainty but for which we otherwise lack constraints from first-principles calculations.
@@
Figure 10: Left panel: Uncertainties in the $q_{T}$ spectrum at N3+1LL relative to the N4LL, due to the nonperturbative parameters (purple band), the TNPs (orange band) and their sum in quadrature band (in dark orange). The different lines show the post-fit relative impact of each nonperturbative parameter in the spectrum. Right panel: Post-fit constraints on the TNPs and nonperturbative parameters (error bars) and their impact on $\alpha_{s}(m_{Z})$, with the solid (dashed) grey band showing the impact of the post-fit downward (upward) variations of each parameter. The stars indicate the true value of the parameters. The post-fit constraints for $\Lambda_{2}$ and $\lambda_{2}$ ($\Lambda_{4}$ and $\lambda_{4}$) are normalized to $0.25\,\mathrm{GeV}^{2}$ ($0.06\,\mathrm{GeV}^{4}$).
@@
In figure 10, we show the result of our nominal fit including nonperturbative parameters. The left panel displays the breakdown of the post-fit uncertainties in the $q_{T}$ spectrum. The uncertainty due to the nonperturbative parameters is shown in violet, the perturbative uncertainty due to TNPs is shown in orange, and their combined uncertainty is represented by the dark orange band, which corresponds to the total post-fit uncertainty for this Asimov test. [^7] When profiling multiple sources of uncertainty, the breakdown of the total uncertainty into individual components is somewhat ambiguous. Here, we obtain the individual bands on the $q_{T}$ spectrum by considering the relevant sub-matrices of the total post-fit covariance matrix. One benefit of the Asimov approach however is that we can also consider different sources in isolation, as was done for the TNP component in section 4.2. We do not show a pre-fit band here, since we do not have pre-fit constraints for $\Lambda_{2,4}$. The perturbative pre-fit uncertainty corresponds to the one shown in figure 6. Overlaid on the plot are the post-fit relative impacts of each nonperturbative parameter. A strong negative correlation emerges between $\lambda_{2}$ and $\Lambda_{2}$, as evident from their significant impact on the spectrum despite the relatively small size of the overall nonperturbative uncertainty. In the right panel of the same figure, we show the constraints on all fit parameters, along with their individual impacts on $\alpha_{s}(m_{Z})$. Compared to the purely perturbative case discussed in section 4.2.3, the post-fit constraints on the TNPs are still consistent with their true values (indicated by stars), but some are now not as strongly constrained by the data. Due to the additional fit parameters introduced, the data has additional freedom now to determine which parameters are most relevant in reducing the uncertainty on the parameter of interest. It is clear that some nonperturbative parameters, especially $\Lambda_{2,4}$, have a large impact on $\alpha_{s}(m_{Z})$.
@@
Figure 11: Uncertainty on $\alpha_{s}(m_{Z})$ from various Asimov tests. Red bars include TNP profiling (i.e., fitting $\alpha_{s}(m_{Z})$ together with all theory nuisance parameters). Below, we show cases excluding TNP profiling, in blue, with different sets of fitted parameters. From top to bottom: fit only $\alpha_{s}(m_{Z})$, fit $\alpha_{s}(m_{Z})$, $\lambda_{2,4}$, $\Lambda_{2,4}$ including lattice QCD constraints for CS kernel parameters, fit $\alpha_{s}(m_{Z})$, $\lambda_{2}$ and $\Lambda_{2,4}$, fit $\alpha_{s}(m_{Z})$ and $\Lambda_{2,4}$. All the uncertainties are in units of $10^{-3}$.
@@
Next, we explore the relationship between the nonperturbative effects and the uncertainty on $\alpha_{s}(m_{Z})$, and how the TNPs behave in this context. Figure 11 presents the results of several Asimov fits as discussed next. For each case, we display two uncertainty bars. The outer red bars show the total uncertainty of the nominal fit simultaneously profiling TNPs and nonperturbative parameters. The inner blue bars show the uncertainty associated with fitting $\alpha_{s}(m_{Z})$ alongside the nonperturbative parameters only. They are obtained by repeating the fit fixing the TNPs to their nominal post-fit central values.
@@
Fortunately, as discussed in section 3.3, the CS kernel can be constrained using lattice QCD data. Incorporating these constraints – via both the central values from Eq. (3.3.2) and the associated covariance matrix in Eq. (3.35) – enables us to include both $\lambda_{2,4}$ in the fit, alongside the TMD PDF parameters $\Lambda_{2,4}$. This yields our nominal fit already shown in figure 10. The $\alpha_{s}(m_{Z})$ uncertainties resulting from this fit are
@@ Appendix A Comments on Recent $\alpha_{s}$ Extraction by ATLAS
In table 2, we summarize our estimates of the expected size of perturbative uncertainties based on fits to Asimov data and compare them to the values quoted in Ref. [33]. For scale variations, we use the quadrature sum of envelopes here. Using instead the naive total envelope ($1.73\times 10^{-3}$) leads to a slightly smaller total perturbative uncertainty of $\pm 2.30\times 10^{-3}$. Our uncertainty estimates are based on fits to Asimov data, so while the precise numbers will differ in a fit to the real data their typical size will be the same. For the uncertainties due to scale variations, we have explicitly checked in preliminary fits to real data that the size of the obtained uncertainty is indeed very similar. We thus have reason to believe that the total theory uncertainty quoted in Ref. [33] is underestimated by a factor of 4 to 5. This would make it by far the dominant uncertainty and as a result also the total uncertainty of $\pm 0.9\times 10^{-3}$ quoted in Ref. [33] would be underestimated by about a factor of 3.
@@
Scale variations $\pm 0.42$ $\pm 2.43$
N4LL approximation $\pm 0.04$
N4LL approximation $\pm 0.75$
Flavor/quark masses +0.40+0.40 0.29-0.29 $\pm 1.32$
Total +0.58+0.58 0.51-0.51 $\pm 2.87$
@@ A.1 Perturbative uncertainties from scale variations
The analysis in Ref. [33] uses the method of enveloped scale variations discussed in section 2.2 to estimate the perturbative theory uncertainty on the resummed $q_{T}$ spectrum. This uncertainty is propagated to the fitted $\alpha_{s}(m_{Z})$ by scanning over individual scale variations. This yields maximum deviations in $\alpha_{s}(m_{Z})$ of $0.23\times 10^{-3}$ and $0.61\times 10^{-3}$, which are symmetrized to the quoted $\pm 0.42\times 10^{-3}$ scale-variation uncertainty on $\alpha_{s}(m_{Z})$.
@@
As we have argued in the main text, extracting $\alpha_{s}(m_{Z})$ from the small-$q_{T}$ spectrum crucially relies on shape effects and is therefore extremely sensitive to the theory uncertainty on the shape of the spectrum and the corresponding bin-by-bin theory correlations. As discussed in detail in section 2.2, these are not correctly captured by scale variations, which means the resulting variations in the fitted $\alpha_{s}(m_{Z})$ cannot be regarded as a trustworthy uncertainty estimate. Scanning over scale variations in section 4.1 we find significantly larger maximum variations in the fitted $\alpha_{s}(m_{Z})$, as seen in the right panel of figure 2. The envelope over all scale variations yields $1.73\times 10^{-3}$, while adding in quadrature the envelopes of different classes of scale variations yields $2.43\times 10^{-3}$, see Eqs. (4.1) and (4.1).
Our scale-variation study is performed at N4LL, which is the currently highest known (essentially) complete perturbative order. It yields uncertainties of $0.5-1\%$ in the range $5\,\mathrm{GeV}<q_{T}<30\,\mathrm{GeV}$ and growing up to $3-4\%$ for $0<q_{T}<5\,\mathrm{GeV}$, as seen in the left panel of figure 2. Unfortunately, from the limited information given in Ref. [33] it is not clear what the exact size of the scale-variation based uncertainties on the $q_{T}$ spectrum are at their highest used order (denoted N4LL in the next subsection). From our own runs of the public version of DYturbo (v1.4.2) we find spectrum uncertainties of $0.5-1\%$ for $q_{T}<30\,\mathrm{GeV}$ and growing to $2-3\%$ for $q_{T}\to 0$. [^10] Similarly, in Ref. [30], numerical results for the $q_{T}$ spectrum at this N4LL order are given for $E_{\mathrm{cm}}=13\,\mathrm{TeV}$ and including fiducial lepton cuts, in which case the spectrum uncertainties are also close to around $1\%$. Therefore, the envelope of scale variations used in Ref. [33] at the level of the $q_{T}$ spectrum should be of similar size to ours here. This means, for similar-size scale variations in the $q_{T}$ spectrum we obtain up to five times larger variations in $\alpha_{s}(m_{Z})$.
@@
One difference which could (at least partially) contribute to this large discrepancy is that the scale-variation recipes used in the SCETlib and DYturbo resummation frameworks differ. Whilst both use the conventional factor of two for varying scales up and down and avoid simultaneous variations that would compound to factors of four in any scale ratios, the specific types of scale variations differ substantially and are more limited in DYturbo. In particular, the low scale $\sim q_{T}\sim 1/b_{T}$ is varied as part of the scale variation setup in SCETlib, while it is kept fixed in DYturbo. We stress that there are no compelling theoretical arguments for keeping this low scale fixed when estimating uncertainties using scale variations. [^11] This low scale is genuine to the resummation problem, as discussed in the following subsection, and therefore must effectively appear in any resummation framework. It is also kept variable in the original CSS framework [75] on which DYturbo is based. In the context of an ongoing benchmarking study of various $q_{T}$ resummation frameworks within the LHC electroweak working group, the need for varying this low scale has been widely recognized by now and, as far as we can tell, has been gradually adopted also by most other resummation frameworks.
@@
As discussed in section 2, in a situation like this where theory correlations are critically important, for the same overall uncertainty in the spectrum, the resulting uncertainty on $\alpha_{s}(m_{Z})$ entirely depends on the specific scale variations that happen to be included in the scanning. A similar breakdown into individual scale variations as in our figure 2 of the results in Ref. [33] might shed some more light on this. Ultimately, however, the large discrepancies we observe between different resummation frameworks using different scale variation recipes likely just reflect the fact that the scale-variation method cannot be relied upon for propagating the perturbative uncertainty from the $q_{T}$ spectrum to $\alpha_{s}(m_{Z})$.
@@
Nevertheless, if scale variations are the method of choice, one must be aware of their limitations and at minimum it must be ensured that a given scale-variation recipe does not underestimate the resulting uncertainty on the parameter of interest. For this purpose, we have clear evidence that the type of scale variations performed in SCETlib yield a more realistic uncertainty for $\alpha_{s}(m_{Z})$. This is because when scanning over TNPs, which do encode the correct theory correlations, spectrum uncertainties of comparable (or even somewhat smaller) size translate into an uncertainty on $\alpha_{s}(m_{Z})$ of $1.75\times 10^{-3}$ (see section 4.2.1 and figure 3), which is similar to what we find with scale variations.
@@
A.2 Different N4LL countings and approximation uncertainties